OpenStudy (anonymous):

For what value of K will the system of equations given below have no solution? 2x + 5y = 9 -3x - K y = 4

OpenStudy (anonymous):

K = 15/2 = 7.5

OpenStudy (anonymous):

given two linear equations in two variables a1x + b1y + c1 = 0 and a2x + b2y + c2 = 0 they will not have a solution if a1 b1 c1 ---- = ---- ╪ ---- a2 b2 c2 Rearranging given equations we get 2x + 5y - 9 = 0 -3x - Ky - 4 = 0 Here obviously a1/a2 ╪ c1/c2 because 2/-3 ╪ -9/-4 so if give value to K such that b1/b2 = a1/a2, then there will be no solution b1 2 so ---- has to be made = ----- b2 -3 5 2 so ----- = ----- -k -3 cross multiplying -15 = -2k -15/-2 = k so k = 15/2 It looks like a long calculation, but once you know what is to be done, you can do it orally and very quickly.......